# arXiv:2602.23780v1[eess.SP] 27 Feb 2026

### 27 Feb 2026 v1

@@ 2.1 Polynomials in the real line
where the fact $\binom{j+1}{k}=\binom{j}{k-1}+\binom{j}{k}$ has been used in the second to last inequality. This proves Equation (6), thus the lemma. ∎
We now prove that the mapping $T_{\varepsilon}$ defined in Corollary 2, is an isomorphism of the vector space $\mathscr{P}_{n}(\mathbb{R})$ onto itself. For this purpose we will use the results from Corollary 1 and Remark 2 in a recursive manner. This result has some remarkable properties: in $\mathscr{P}_{n}(\mathbb{R})$ the convolution with an even convolution Schwartz kernel is always a reversible operation. Therefore, we can consider this to be an extension of the Weierstrass Transform of polynomials, in which $\varphi$ is the Gaussian Kernel, and so $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ is even. See [bilodeau1962weierstrass] for a discussion of Hermite polynomials and the Weierstrass function. Finally, note in the proof of the following Theorem that the even part plays a crucial role. Indeed, the even part accounts for two reasons. First, if $p\in\mathscr{P}_{n}(\mathbb{R})$ then $T_{\varepsilon}(p)=p+p_{1}$ with $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$, that is, it reduces $p_{1}$ two degrees below $p$. Also if $p$ is an affine or constant function then $T_{\varepsilon}(p)=p$. This is what will allow us to prove the automorphism and the computation of the inverse.
Theorem 2.
Let $n\in\mathbb{N}$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Then $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$, defined as in Corollary 2 is an automorphism. Moreover, given $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n$ even
@@
Proof.
The fact that $T_{\varepsilon}$ is linear follows directly from the definition of the convolution. We proceed to prove that it is bijective. Suppose first that $n=2$. We know by Corollary 2 that $T_{\varepsilon}(p)=p+p_{1}$ where $p_{1}\in\mathscr{P}_{0}(\mathbb{R})$, and therefore $T_{\varepsilon}(p_{1})=p_{1}$. If $p,q\in\mathscr{P}_{2}(\mathbb{R})$ such that $p(x)=a_{0}+a_{1}x+a_{2}x^{2}$ and $q(x)=b_{0}+b_{1}x+b_{2}x^{2}$ for all $x\in\mathbb{R}$, then $p_{1}(x)=a_{2}\varepsilon^{2}c_{2}$ and $q_{1}(x)=b_{2}\varepsilon^{2}c_{2}$. Thus
@@
but since $p+p_{1}\in\mathscr{P}_{2}(\mathbb{R})$ and $q+q_{1}\in\mathscr{P}_{2}(\mathbb{R})$ with the same leading coefficients as $p$ and $q$ by Corollary 1, and two polynomials of the same degree are equal if and only if all of its coefficients are equal, it follows that
@@
  • $T_{\varepsilon}$ is injective. Let $n\in\mathbb{N}$. Suppose $T(p)=T(q)$, then $p+p_{1}=q+q_{1}$. Since $p_{1},q_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ this implies that $a_{n}=b_{n}$ and $a_{n-1}=b_{n-1}$. Now, note that by Remark 2 that
@@
proving the first equality of Equation (7). The second equality comes from Lemma 1. Finally, the third equality can we proven by induction. For $n=2$ we have that $T_{\varepsilon}^{-1}(p)=2p-T_{\varepsilon}(p)$ for both formulas. Suppose then it holds for $n\geq 2$ with $n$ even. Then for $n+2$
@@
$=\sum_{j=0}^{n/2}(-1)^{\frac{n}{2}-j}\binom{\frac{n}{2}+1}{j}T_{\varepsilon}^{\frac{n}{2}-j}(p)+\sum_{k=1}^{\frac{n}{2}+1}(-1)^{\frac{n}{2}+1-k}\binom{\frac{n}{2}+1}{k}T_{\varepsilon}^{\frac{n}{2}+1-k}(p)$
$+(-1)^{\frac{n}{2}+1}T_{\varepsilon}^{\frac{n}{2}+1}(p)$
@@
proving Equation (7). The proof for the case where $n$ odd, i.e., Equation (8), is identical to the even case. ∎
Corollary 3.
@@
Proof.
The first result follows from a change of indexes in the sum. The second follows noting that, if $f:V\to V$ is a linear mapping, where $V$ is any vector space, then, by the binomial theorem for linear functions, $\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}f^{k}=(\operatorname{id}-f)^{n}$ where $\operatorname{id}:V\to V$ is the identity on $V$. ∎
Note from Corollary 3 it is only needed to compute $n/2$ convolutions numerically speaking, because each time the index of the sum advances, it is only needed to convolve the previously convolved function.
Remark 3.
Theorem 2 states that, for polynomials, convolution is invertible as long as the kernel is an even (Schwartz) function. Note that convolution is, practically speaking, the filtering of the function with a smooth filter in $\mathscr{S}_{M}(\mathbb{R})$. Note that since $\mathscr{S}_{M}(\mathbb{R})\subset\mathscr{S}(\mathbb{R})\subset L^{2}(\mathbb{R})$ the filter has finite energy. What the Theorem states is that, even after having (for example) low-filtered a polynomial via convolution with an even convolution kernel (e.g., Gaussian Kernel, or the compactly supported convolution kernel $x\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$), we can recover the original function uniquely. This powerful fact can also be numerically implemented easily by Equations (7) and (8). See Section 3 for an application related to this remark.
Remark 4.
@@ 2.2 The extension to a bigger space of functions: The deconvolution of arbitrary functions
The natural question now is to ask if, for a given function $f$, it holds (under certain assumptions about $f$) that, by passing to the limit in Equation (7), we can recover $f$ from $T_{\varepsilon}(f)$. That is, if $T_{\varepsilon,n}^{-1}$ is the inverse of the automorphism $T_{\varepsilon}$ in $\mathscr{P}_{n}(\mathbb{R})$ for some even $\varphi\in\mathscr{S}(\mathbb{R})$ (see Theorem 2) does it hold that
@@
in some sense of convergence? The answer is affirmative under some conditions on the Fourier transform of $\varphi\in\mathscr{S}$ and different assumptions for $f$. We prove these results in the next theorem, in which the assumption that $\varphi$ is even is not always required. As mention in the introduction, we will denote $\mathcal{F}\left(f\right)$ or $\widehat{f}$ the Fourier transform of a function $f$ whenever it exists, and as $\mathcal{F}$ the Fourier transform as an operator. For its properties refer to [grubb2008distributions] and [grafakos2008classical].
Theorem 3.
@@
    First note that since $f\in L^{2}(\mathbb{R})$ and $\varphi_{\varepsilon}\in\mathscr{S}(\mathbb{R})$ then $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}\in L^{2}(\mathbb{R})$ by Young’s inequality for convolution [folland1999real], which implies that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\in L^{2}(\mathbb{R})$ and so the operator makes sense. By the Fourier transform properties, it holds that $\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)\in L^{2}(\mathbb{R})$ for all $n\in\mathbb{N}$ and all $f\in L^{2}(\mathbb{R})$. Moreover $\widehat{\varphi_{\varepsilon}}\in\mathscr{S}(\mathbb{R})$, and since $\mathcal{F}\left(T_{\varepsilon}(f)\right)=\widehat{f}\widehat{\varphi_{\varepsilon}}$, by induction it is easily seen that $\mathcal{F}\left(T^{k}_{\varepsilon}(f)\right)=\widehat{f}(\widehat{\varphi_{\varepsilon}}\dots\widehat{\varphi_{\varepsilon}})=\widehat{f}\widehat{\varphi_{\varepsilon}}^{k}$ for all $k\in\mathbb{N}$. Thus, by the linearity of the Fourier transform
@@
    $=\lim_{n\to\infty}\mathcal{F}^{-1}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)),$
@@
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))},\psi\rangle$ $=\langle\delta\cdot(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}),\psi\rangle$
    $=\langle\delta,(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle=(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})(0)\psi(0).$
@@
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathrm{pv}\frac{1}{x}))},\psi\rangle$ $=\langle-i\pi\operatorname{sign},(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle$
    $=-i\pi\int_{\mathbb{R}}\operatorname{sign}(\xi)(1-(1-\widehat{\varphi_{\varepsilon}}(\xi))^{n+1})\psi(\xi)d\xi,$
@@
    $=\lim_{n\to\infty}\mathcal{F}^{-1}\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)=\lim_{n\to\infty}T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)),$
    in the Schwartz topology. If $f\in\mathscr{S}^{\prime}(\mathbb{R})$, since $\varphi$ is even,
@@
    So we proceed now to prove that for any $\psi\in\mathscr{S}(\mathbb{R})$ then $\psi(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ as $n\to\infty$ in the Schwartz topology if $\varphi$ is the Gaussian Kernel [^2] Note we cannot prove that $1-(1-\widehat{\varphi_{\varepsilon}})^{n}\to 1$ in the Schwartz topology since the constant function $1$ is not in $\mathscr{S}(\mathbb{R})$, and $\mathscr{S}(\mathbb{R})$ is complete.. Assume without loss of generality that $\varepsilon=1$. Then $\widehat{\varphi}(\xi)=e^{-\xi^{2}}$ for all $\xi\in\mathbb{R}$. Let $h(x)=1-e^{-x^{2}}$ for all $x\in\mathbb{R}$. We need to prove that for all $\alpha,\beta\in\mathbb{N}_{0}$,
@@
    Fix $\alpha,\beta\in\mathbb{N}_{0}$. For simplicity of notation, we will denote both the function $f$ and its value at $x$ as $f(x)$. First, it is not difficult to check by induction that for any $m\in\mathbb{N}$, it holds $D^{m}h(x)=P(m,x)e^{-x^{2}}$ where $P(m,x)$ is a polynomial of degree up to $m$. Applying Faa di Bruno’s formula [roman1980formula] for $m\leq n$, it holds that
@@
    $\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta}(\psi h^{n})(x)\right|$
    $\leq\sum_{m=0}^{\beta}\sum\binom{\beta}{m}\pi(m)\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta-m}\psi(x)\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}P(m!,x)e^{-\bar{k}x^{2}}\right|,$
@@
We want to remark that the last statement of Theorem 3 is a new formula for the inverse of the Weierstrass Transform, i.e., the operator $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}$ is the inverse for the Weierstrass Transform with parameter $\varepsilon>0$ for any $f\in\mathscr{S}(\mathbb{R})$ or $f\in\mathscr{S}^{\prime}(\mathbb{R})$. Note that $L^{p}(\mathbb{R})$ is injected in $\mathscr{S}^{\prime}(\mathbb{R})$ for any $p\in[1,\infty]$, as well as any $v\in L^{1}_{loc}(\mathbb{R})$ with $|v(x)|\leq C(1+|x|^{2})^{N/2}$ for some $N\in\mathbb{N}$ and $C>0$. Also note that the Delta, Heaviside and Principal Value distributions are in $\mathscr{S}^{\prime}(\mathbb{R})$ and so any distribution with compact support. Therefore, Theorem 3 presents a result that can be used in several situations which is numerically simple to compute. To end this section, note that if $f\in\mathscr{S}(\mathbb{R})$ of $f\in\mathscr{S}^{\prime}(\mathbb{R})$ and $\varphi$ is the Gaussian Kernel, then, in their respective topologies
@@
In Section 3 we show the numerical validation of Theorem 3.

### 17 Apr 2026 v2

@@ 2.1 Polynomials in the real line
where the fact $\binom{j+1}{k}=\binom{j}{k-1}+\binom{j}{k}$ has been used in the second to last inequality. This proves Equation (6), thus the lemma.
We now prove that the mapping $T_{\varepsilon}$, defined in Corollary 2.8, is an isomorphism on the vector space $\mathscr{P}_{n}(\mathbb{R})$. Rather than exploiting the finite-dimensional nature of $\mathscr{P}_{n}(\mathbb{R})$ to construct the matrix for the linear isomorphism $T_{\varepsilon}$, our proof proceeds recursively by using the results established in Corollary 2.5 and Remark 2.7. We adopt this recursive approach because it serves as the foundational mechanism that will allow us to extend the inversion to infinite-dimensional spaces of arbitrary functions and distributions in Theorem 2.18. Nevertheless, this isomorphism establishes a remarkable structural property: within $\mathscr{P}_{n}(\mathbb{R})$, convolution with an even Schwartz kernel is always an exactly reversible operation. Consequently, this framework can be viewed as an extension of the Weierstrass transform applied to polynomials, where $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ represents the even Gaussian kernel (see bilodeau1962weierstrass for a related discussion on Hermite polynomials and the Weierstrass function). Finally, we emphasize that the parity of the convolution kernel plays a critical role in the following proof. Its necessity is twofold: first, for any $p\in\mathscr{P}_{n}(\mathbb{R})$, the operator yields $T_{\varepsilon}(p)=p+p_{1}$, where $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$. Second, it ensures that affine and constant functions remain completely invariant (i.e., $T_{\varepsilon}(p)=p$). This degree reduction is precisely what enables both the proof of the automorphism and the explicit, finite-sum computation of its inverse.
Theorem 2.12.
Let $n\in\mathbb{N}$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Then $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$, defined as in Corollary 2.8 is an automorphism. Moreover, given $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n$ even
@@
Proof 2.13.
The fact that $T_{\varepsilon}$ is linear follows directly from the definition of the convolution. We proceed to prove that it is bijective. Suppose first that $n=2$. We know by Corollary 2.8 that $T_{\varepsilon}(p)=p+p_{1}$ where $p_{1}\in\mathscr{P}_{0}(\mathbb{R})$, and therefore $T_{\varepsilon}(p_{1})=p_{1}$. If $p,q\in\mathscr{P}_{2}(\mathbb{R})$ such that $p(x)=a_{0}+a_{1}x+a_{2}x^{2}$ and $q(x)=b_{0}+b_{1}x+b_{2}x^{2}$ for all $x\in\mathbb{R}$, then $p_{1}(x)=a_{2}\varepsilon^{2}c_{2}$ and $q_{1}(x)=b_{2}\varepsilon^{2}c_{2}$. Thus
@@
but since $p+p_{1}\in\mathscr{P}_{2}(\mathbb{R})$ and $q+q_{1}\in\mathscr{P}_{2}(\mathbb{R})$ with the same leading coefficients as $p$ and $q$ by Corollary 2.5, and two polynomials of the same degree are equal if and only if all of its coefficients are equal, it follows that
@@
  • $T_{\varepsilon}$ is injective. Let $n\in\mathbb{N}$. Suppose $T(p)=T(q)$, then $p+p_{1}=q+q_{1}$. Since $p_{1},q_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ this implies that $a_{n}=b_{n}$ and $a_{n-1}=b_{n-1}$. Now, note that by Remark 2.7 that
@@
proving the first equality of Equation (7). The second equality comes from Lemma 2.10. Finally, the third equality can we proven by induction. For $n=2$ we have that $T_{\varepsilon}^{-1}(p)=2p-T_{\varepsilon}(p)$ for both formulas. Suppose then it holds for $n\geq 2$ with $n$ even. Then for $n+2$
@@
$=\sum_{j=0}^{n/2}(-1)^{\frac{n}{2}-j}\binom{\frac{n}{2}+1}{j}T_{\varepsilon}^{\frac{n}{2}-j}(p)+\sum_{k=1}^{\frac{n}{2}+1}(-1)^{\frac{n}{2}+1-k}\binom{\frac{n}{2}+1}{k}T_{\varepsilon}^{\frac{n}{2}+1-k}(p)+(-1)^{\frac{n}{2}+1}T_{\varepsilon}^{\frac{n}{2}+1}(p)$
@@
proving Equation (7). The proof for the case where $n$ odd, i.e., Equation (8), is identical to the even case.
Corollary 2.14.
@@
Proof 2.15.
The first result follows from a change of indexes in the sum. The second follows noting that, if $f:V\to V$ is a linear mapping, where $V$ is any vector space, then, by the binomial theorem for linear functions, $\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}f^{k}=(\operatorname{id}-f)^{n}$ where $\operatorname{id}:V\to V$ is the identity on $V$.
Note from Corollary 2.14 it is only needed to compute $n/2$ convolutions numerically speaking, because each time the index of the sum advances, it is only needed to convolve the previously convolved function.
Remark 2.16.
Theorem 2.12 states that, for polynomials, convolution is invertible as long as the kernel is an even (Schwartz) function. Note that convolution is, practically speaking, the filtering of the function with a smooth filter in $\mathscr{S}_{M}(\mathbb{R})$. Note that since $\mathscr{S}_{M}(\mathbb{R})\subset\mathscr{S}(\mathbb{R})\subset L^{2}(\mathbb{R})$ the filter has finite energy. What the Theorem states is that, even after having (for example) low-filtered a polynomial via convolution with an even convolution kernel (e.g., Gaussian Kernel, or the compactly supported convolution kernel $x\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$), we can recover the original function uniquely. This powerful fact can also be numerically implemented easily by Equations (7) and (8). See Section 3 for an application related to this remark.
Remark 2.17.
@@ 2.2 The extension to a bigger space of functions: the deconvolution of arbitrary functions
The natural question now is to ask if, for a given function $f$, it holds (under certain assumptions about $f$) that, by passing to the limit in Equation (7), we can recover $f$ from $T_{\varepsilon}(f)$. That is, if $T_{\varepsilon,n}^{-1}$ is the inverse of the automorphism $T_{\varepsilon}$ in $\mathscr{P}_{n}(\mathbb{R})$ for some even $\varphi\in\mathscr{S}(\mathbb{R})$ (see Theorem 2.12) does it hold that
@@
in some sense of convergence? The answer is affirmative under some conditions on the Fourier transform of $\varphi\in\mathscr{S}$ and different assumptions for $f$. We prove these results in the next theorem, in which the assumption that $\varphi$ is even is not always required. As mention in the introduction, we will denote $\mathcal{F}\left(f\right)$ or $\widehat{f}$ the Fourier transform of a function $f$ whenever it exists, and as $\mathcal{F}$ the Fourier transform as an operator. For its properties refer to grubb2008distributions and grafakos2008classical.
Theorem 2.18.
@@
    First note that since $f\in L^{2}(\mathbb{R})$ and $\varphi_{\varepsilon}\in\mathscr{S}(\mathbb{R})$ then $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}\in L^{2}(\mathbb{R})$ by Young’s inequality for convolution folland1999real, which implies that $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\in L^{2}(\mathbb{R})$ and so the operator makes sense. By the Fourier transform properties, it holds that $\mathcal{F}\left(T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))\right)\in L^{2}(\mathbb{R})$ for all $n\in\mathbb{N}$ and all $f\in L^{2}(\mathbb{R})$. Moreover $\widehat{\varphi_{\varepsilon}}\in\mathscr{S}(\mathbb{R})$, and since $\mathcal{F}\left(T_{\varepsilon}(f)\right)=\widehat{f}\widehat{\varphi_{\varepsilon}}$, by induction it is easily seen that $\mathcal{F}\left(T^{k}_{\varepsilon}(f)\right)=\widehat{f}(\widehat{\varphi_{\varepsilon}}\cdots\widehat{\varphi_{\varepsilon}})=\widehat{f}\widehat{\varphi_{\varepsilon}}^{k}$ for all $k\in\mathbb{N}$. Thus, by the linearity of the Fourier transform
@@
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathds{1}))},\psi\rangle$ $=\langle\delta\cdot(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1}),\psi\rangle=\langle\delta,(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle=(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})(0)\psi(0).$
@@
    $\langle\widehat{T_{\varepsilon,n}^{-1}(T_{\varepsilon}(\mathrm{pv}\frac{1}{x}))},\psi\rangle$ $=\langle-i\pi\operatorname{sign},(1-(1-\widehat{\varphi_{\varepsilon}})^{n+1})\psi\rangle=-i\pi\int_{\mathbb{R}}\operatorname{sign}(\xi)(1-(1-\widehat{\varphi_{\varepsilon}}(\xi))^{n+1})\psi(\xi)d\xi,$
@@
    in the Schwartz topology. If $f\in\mathscr{S}^{\prime}(\mathbb{R})$, being $\varphi$ even,
@@
    So we proceed now to prove that for any $\psi\in\mathscr{S}(\mathbb{R})$ then $\psi(1-\widehat{\varphi_{\varepsilon}})^{n}\to 0$ as $n\to\infty$ in the Schwartz topology if $\varphi$ is the Gaussian Kernel. Assume without loss of generality that $\varepsilon=1$. Then $\widehat{\varphi}(\xi)=e^{-\xi^{2}}$ for all $\xi\in\mathbb{R}$. Let $h(x)=1-e^{-x^{2}}$ for all $x\in\mathbb{R}$. We need to prove that for all $\alpha,\beta\in\mathbb{N}_{0}$,
@@
    Fix $\alpha,\beta\in\mathbb{N}_{0}$. For simplicity of notation, we will denote both the function $f$ and its value at $x$ as $f(x)$. First, it is not difficult to check by induction that for any $m\in\mathbb{N}$, it holds $D^{m}h(x)=P(m,x)e^{-x^{2}}$ where $P(m,x)$ is a polynomial of degree up to $m$. Applying Faa di Bruno’s formula roman1980formula for $m\leq n$, it holds that
@@
    $\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta}(\psi h^{n})(x)\right|\leq\sum_{m=0}^{\beta}\sum\binom{\beta}{m}\pi(m)\sup_{x\in\mathbb{R}}\left|x^{\alpha}D^{\beta-m}\psi(x)\frac{n!}{(n-\bar{k})!}(1-e^{-x^{2}})^{n-\bar{k}}P(m!,x)e^{-\bar{k}x^{2}}\right|,$
@@
We want to remark that the last statement of Theorem 2.18 is a new formula for the inverse of the Weierstrass Transform, i.e., the operator $\lim_{n\to\infty}T_{\varepsilon,n}^{-1}$ is the inverse for the Weierstrass Transform with parameter $\varepsilon>0$ for any $f\in\mathscr{S}(\mathbb{R})$ or $f\in\mathscr{S}^{\prime}(\mathbb{R})$. Note that $L^{p}(\mathbb{R})$ is injected in $\mathscr{S}^{\prime}(\mathbb{R})$ for any $p\in[1,\infty]$, as well the set $\{v\in L^{1}_{loc}(\mathbb{R})\mid|v(x)|\leq C(1+|x|^{2})^{N/2}\text{ for some $N\in\mathbb{N}$ and $C>0$}\}$. Also note that the Delta, Heaviside and Principal Value distributions are in $\mathscr{S}^{\prime}(\mathbb{R})$ and so any distribution with compact support. Therefore, Theorem 2.18 presents a result that can be used in several situations which is numerically simple to compute. To end this section, note that if $f\in\mathscr{S}(\mathbb{R})$ of $f\in\mathscr{S}^{\prime}(\mathbb{R})$ and $\varphi$ is the Gaussian Kernel, then, in their respective topologies
@@
In Section 3 we show the numerical validation of Theorem 2.18.